Down-To-Earth sells houses. The following table represents the selling price of a house (y) in thousands of rands and the number of houses sold at that price (x):

x 5 15 19 7

y 500 900 1500 2 000

The standard deviation for the number of houses sold is

[1] 6,6.

[2] 4,0.

[3] 11,5.

[4] 5,7.

[5] none of the above.



1
Expert's answer
2021-07-19T10:33:22-0400

we have to find the standard deviation for the number of

 houses sold

so  the given data is 5,15,19,7

mean "\bar{x}=\frac{5+15+19+7}{5}=11.5"

standard deviation (s)"=\sqrt{\frac{\displaystyle\sum_{1}(x_1-\bar{x})}{n-1}}"

"=\frac{\sqrt{(5-11.5 )^2+(15-11.5)^2 +(19-11.5)^2 +(7-11.5)^2}}{5-1}"

"=\sqrt{43.666667}\\=6.6"